The equation of motion of classical harmonic oscillator is expressed by x=asin(omega)t.Show that probability of finding the particle between x and x+dx is given by,
P(x) dx=(dx)÷ (π√[d^2-x^2])
1
Expert's answer
2020-03-30T08:08:26-0400

As per the question,

wave equation is given as "x=a\sin(\omega)t"

Let "dP=\dfrac{2dt}{T}-------------(i)"

Where T is the time period and dt is the time taken in x to x+dx

"x=a\sin(\omega)t"

"dx=a\omega \cos(\omega t)dt"

"\Rightarrow dt=\dfrac{dx}{a\omega \cos(\omega t)}"

now

from the equation (i),

"dP=\dfrac{2dx}{a\omega\cos(\omega t)T}"

we know that "\omega=\dfrac{2\pi}{T}"

"\Rightarrow dP=\dfrac{2dx}{a\dfrac{2\pi}{T}\cos(\omega t)T}=\dfrac{2dx}{2\pi a\cos(\omega t)}"

now we know that "x=a\sin(\omega)t"

"\sin(\omega t)=\dfrac{x}{a}"

"\cos(\omega t)=\sqrt{1-\sin^2(\omega t)}"

"\Rightarrow \cos(\omega t)=\sqrt{1-(\dfrac{x}{a})^2}"

"\Rightarrow dP=\dfrac{2dx}{2\pi a\cos(\omega t)}"

"\Rightarrow dP=\dfrac{2dx}{2\pi a\sqrt{1-(\dfrac{x}{a})^2}}"

"\Rightarrow dP=\dfrac{dx}{\pi \sqrt{a^2-(x)^2}}"

"\Rightarrow \dfrac{dP}{dx}=\dfrac{1}{\pi \sqrt{a^2-x^2}}"